# NLTE spectral modelling for a carbon-oxygen and helium white-dwarf merger as a Ca-rich transient candidate

### v1

@@
The progenitors of Ca-rich transients are still unknown. However, the fact that they are typically found in remote locations in old host environments (Perets et al. 2010; Foley 2015; Lunnan et al. 2017) suggests older progenitors that involve white dwarfs (WDs) in binary systems. Various models have been proposed to explain Ca-rich transients (see Shen et al. 2019 for a review).
@@ 2.1 NLTE and non-thermal radiative transfer
The radiative transfer simulations are carried out using the time-dependent Monte Carlo radiative transfer code artis222https://github.com/artis-mcrt/artis/ (Sim 2007; Kromer & Sim 2009; Bulla et al. 2015; Shingles et al. 2020). artis follows the methods of Lucy (2002, 2003, 2005) and is based on dividing the radiation field into indivisible energy packet Monte Carlo quanta (packets hereafter). Here, we utilise the full NLTE and non-thermal capabilities added to artis by Shingles et al. (2020). This includes a NLTE population and ionization solver and treatment for collisions with non-thermal leptons. These improvements both extend the validity of our radiative transfer simulations to the nebular phase and allow the spectral contribution of He to be predicted. To follow the energy distribution of high-energy leptons, which result from nuclear decays and Compton scattering of -rays, artis solves the Spencer-Fano equation (as framed by Kozma & Fransson 1992). Auger electrons are allowed to contribute to heating, ionization and excitation of bound electrons by non-thermal collisions is also included. artis utilises the full Monte Carlo photon-packet trajectories to obtain a rate estimator for each photoionization cross-section included while a parameterized radiation field is used to estimate bound-bound transition rates. The atomic data set we use is based on the compilation of cmfgen (Hillier 1990; Hillier & Miller 1998) and is similar to that described by Shingles et al. (2020) but with some additional species and ions including atomic data for Ca IV, Ti II and Ti III from the most recent compilation of cmfgen (see Blondin et al. 2023). In this simulation we include He i-iii, C i-iv, O i-iv, Na i-iii, Ne i-iii, Mg i-iii, Al i-iv Si i-iv, S i-iv, Ar i-iv, Ca i-v, Ti ii-iv, Cr i-v, Fe i-v, Co ii-v and Ni ii-v. We note that there is no Ti i data in our atomic data set. We do not expect Ti i to become the dominant ion in our simulations (we do include both Ca i and Cr i, which remain subdominant ions at all times). However, it is possible that Ti i could contribute to the spectrum in the later phases we consider when the temperature and ionization state are low. However, we do not expect that this would significantly impact our principle conclusions on the formation of He i and Ca ii features.
@@
We initialised the nebular phase simulation at 40 d post explosion and followed the spectroscopic evolution using packets until 70 d post explosion with logarithmically spaced time steps (). The initial time steps of the nebular phase simulation were also treated in LTE and the NLTE solver was switched on at time step 12 (44 d).
@@ 2.2 CO He WD merger ejecta model
While artis is capable of multi-dimensional simulations, owing to the increased computational cost of including our full NLTE treatment, we work in 1D for this first exploratory study and employ a spherically averaged model. The full 3D model has clear asymmetries (see Morán-Fraile et al. 2024 or Kozyreva et al. 2024) which likely impact the strengths and shapes of spectral features. However, the 1D average preserves the main characteristics of the model which allows us to address the key questions of whether He i and [Ca ii] line formation are plausible. We construct our 1D model by spherically averaging the 3D hydrodynamic simulation into 87 radial bins. The composition of this 1D ejecta model is shown in Figure 1.
@@ 3 Results
Here we present the optical and NIR spectroscopic evolution of our simulations from peak until the nebular phase333The spectra presented here will be made available on the Heidelberg supernova model archive HESMA (Kromer et al. 2017, https://hesma.h-its.org)..
@@ 3.1 Optical spectroscopic evolution
Ti ii remains a key species in the spectrum, contributing significantly across a range of wavelengths spanning the entire optical spectrum. There is a significant increase in the contributions of neutral species such as Fe i, Ca i, Cr i and Si i. This is a result of the reduced ionization state of the simulation at this epoch. In particular, Fe i has a substantial contribution across a wide range of optical wavelengths and the Fe i feature predicted at Å is the most prominent feature in the spectrum other than the Ca ii/[Ca ii] features discussed above. The complex of emission features predicted by the simulation between - Å is also primarily attributed to the blended contributions of neutral species (Fe i, Ca i, Cr i).
@@ 3.2 NIR spectroscopic evolution
The most prominent feature in the peak NIR spectrum predicted by our simulation is the strong He i 10830 Å absorption feature (see Figure 3), while there is also a clear He i 20587 Å absorption feature. We note that substantial ejected He masses, such as the ejected for our model, are likely required for a He i 20587 Å feature to appear: previous artis simulations of CO WD explosions, ejecting of He, predicted clear He i 10830 Å features, but no obvious He i 20587 Å feature formed (Collins et al. 2023; Callan et al. 2024). A prominent Ca ii 11839, 11950 Å absorption feature is also present at Å. The NIR spectrum is dominated by Ti ii emission for wavelengths shorter than 16000 Å while redward of this He i provides the greatest contribution to the spectrum.
@@ 4.2 Model comparisons with observations in nebular phase
As can be seen from Figure 5 the nebular phase spectrum predicted by our simulation is dominated by the strong [Ca ii] and Ca ii emission features at around 7300 and 8700Å respectively. The simulated features have comparable strengths and velocities to the features observed in the nebular phase spectra of Ca-rich transients. However, our simulation predicts the Ca ii NIR triplet emission to be stronger than the [Ca ii] 7291, 7324 Å emission while these features have roughly equivalent strengths in the observed spectra. Additionally, while we note the observed Ca ii emission features show clear variations in their widths and velocities between different Ca-rich transients (see Figure 5), our simulation predicts Ca ii features that are too broad relative to the observations.
@@
Redward of 7000 Å the observed Ca-rich transients show very substantial variation in their nebular phase spectra and as such the model has varying success in matching the observed spectra over this wavelength range. One feature that appears to be common to all the observations that we compare with here is the emission feature around 5500 Å, which the model attributes primarily to Fe i emission. Other than SN 2022oqm, all the Ca-rich transients which we compare to predict an emission feature around 6300 Å attributed to [O i] 6300, 6364 Å. However, while the simulation predicts such an emission feature the contribution of [O i] 6300, 6364 Å is quite minimal with blended emission from a variety of species including Fe i, Cr i, Ca i driving the formation of the feature (see Figure 2). While this suggests the [O i] 6300, 6364 Å feature is not a clean diagnostic at this epoch it will become more reliable at later times as forbidden emission increasingly dominates over the contribution of allowed transitions to the spectrum.
@@ 5 Conclusions
Therefore we conclude that low-mass CO WDHe WD mergers are promising candidates for Ca-rich transients. However, it is not yet clear whether such models can account for the full range of observed objects in this class of transients. Our model shows a strong spectral contribution from Ti ii at all epochs resulting in a SED at peak that is substantially redder than the majority of Ca-rich transients. The He i 6678, 7065 Å absorption features predicted at peak by our simulation also appear to be more prominent than in the majority of Ca-rich transients. The strong [Ca ii]/Ca ii emission features that dominate our simulated nebular spectra are also too broad relative to observations.
Future work should therefore investigate whether these discrepancies can be resolved. In particular, the He and Ti distributions vary significantly with line-of-sight in the 3D hydrodynamic merger simulation, suggesting that observer orientation may be important. The strength of He and Ti features can also be impacted by factors relevant to their ionization state such as the distribution of in the ejecta which also shows significant variation with viewing angle in the 3D ejecta model. Additionally, the [Ca ii] feature can show asymmetric profiles and in some cases blue-shifts for observed Ca-rich transients, which may indicate that a multi-dimensional structure is present. Future multi-dimensional NLTE radiative transfer simulations should therefore be carried out to investigate if variation between lines-of-sight in the 3D ejecta model can better capture the variety observed for the Ca-rich class.
Ca-rich transients show a variation of approximately three magnitudes in peak r-band brightness. Therefore, further hydrodynamic merger simulations should be carried out to explore if systems with different combinations of CO and He WD masses can robustly explode resulting in a diversity in peak brightness which can cover the variation observed for the Ca-rich class. Additionally, if systems with a lower mass He WD can explode, this should result in a reduction in the He and Ti ejected in the explosion. Future simultaneous optical and NIR observations (e.g. from SOXS) will be ideal for the testing of such models by capturing both the overall SED and the key He i NIR features.
@@ Acknowledgements.
FPC and SAS would like to acknowledge support from the UK Science and Technology Facilities Council (STFC, grant number ST/X00094X/1). This work used the DiRAC Memory Intensive service (Cosma8) at Durham University, managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Durham was funded by BEIS, UKRI and STFC capital funding, Durham University and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). AH is a fellow of the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD) and acknowledges financial support from IMPRS-HD. This work has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101152610. LJS acknowledges support by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (ERC Advanced Grant KILONOVA No. 885281). JMP acknowledges the support of the Department for Economy (DfE). This work received support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 759253 and 945806, the Klaus Tschira Foundation, and the High Performance and Cloud Computing Group at the Zentrum für Datenverarbeitung of the University of Tübingen, the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant no INST 37/935-1 FUGG. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – RO 3676/7-1, project number 537700965, and by the European Union (ERC, ExCEED, project number 101096243). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. SAS thanks the Cosmic Dawn Center / Niels Bohr Institute (University of Copenhagen) who hosted his sabbatical visit during which part of this work was carried out. NumPy and SciPy (Oliphant 2007), Matplotlib (Hunter 2007) and artistools444https://github.com/artis-mcrt/artistools/ (Shingles et al. 2025) were used for data processing and plotting.References
@@
- Callan et al. (2024) Callan, F. P., Collins, C. E., Sim, S. A., et al. 2024, arXiv e-prints, arXiv:2408.03048

### v2

@@
The progenitors of Ca-rich transients are still unknown. However, the fact that they are typically found in remote locations in old host environments (Perets et al. 2010; Foley 2015; Lunnan et al. 2017) suggests older progenitors that involve white dwarfs (WDs) in binary systems. Various models have been proposed to explain Ca-rich transients including He shell detonations (e.g. Perets et al. 2010; Shen et al. 2010; Waldman et al. 2011; Woosley & Kasen 2011), tidal disruptions of white dwarfs either in a compact object binary (e.g. Margalit & Metzger 2016; Zenati et al. 2023) or by an intermediate mass black hole (e.g. Rosswog et al. 2009; Sell et al. 2015) and ultrastipped core-collapse supernovae (e.g. Kawabata et al. 2010) [see e.g. Shen et al. 2019 for a review.]
@@ 2.1 NLTE and non-thermal radiative transfer
The radiative transfer simulations are carried out using the time-dependent Monte Carlo radiative transfer code artis222https://github.com/artis-mcrt/artis/ (Sim 2007; Kromer & Sim 2009; Bulla et al. 2015; Shingles et al. 2020). artis follows the methods of Lucy (2002, 2003, 2005) and is based on dividing the radiation field into indivisible energy packet Monte Carlo quanta (packets hereafter). Here, we utilise the full NLTE and non-thermal capabilities added to artis by Shingles et al. (2020). This includes a NLTE population and ionization solver and treatment for collisions with non-thermal leptons. These improvements both extend the validity of our radiative transfer simulations to the nebular phase and allow the spectral contribution of He to be predicted. To follow the energy distribution of high-energy leptons, which result from nuclear decays and Compton scattering of -rays, artis solves the Spencer-Fano equation (as framed by Kozma & Fransson 1992). Auger electrons are allowed to contribute to heating, ionization and excitation of bound electrons by non-thermal collisions is also included. artis utilises the full Monte Carlo photon-packet trajectories to obtain a rate estimator for each photoionization cross-section included while a parameterized radiation field is used to estimate bound-bound transition rates. The atomic data sets we use are based on the compilation of cmfgen (Hillier 1990; Hillier & Miller 1998) and are similar to that described by Shingles et al. (2020) but with some additional species and ions including atomic data for Ca IV, Ti II, Ti III, Co i and Ni i from the most recent compilation of cmfgen (see Blondin et al. 2023). In our photospheric phase simulation we include He i-iii, C i-iv, O i-iv, Na i-iii, Ne i-iii, Mg i-iii, Al i-iv Si i-iv, S i-iv, Ar i-iv, Ca i-v, Ti ii-iv, Cr i-v, Fe i-v, Co ii-v and Ni ii-v while we also include Co i and Ni i in our nebular phase simulation. We note that there is no Ti i data in our atomic data set. We do not expect Ti i to become the dominant ion in our simulations (we do include both Ca i and Cr i, which remain subdominant ions at all times). However, it is possible that Ti i could contribute to the spectrum in the later phases we consider when the temperature and ionization state are low. However, we do not expect that this would significantly impact our principle conclusions on the formation of He i and Ca ii features.
@@
We initialised the nebular phase simulation at 25 d post explosion and followed the spectroscopic evolution using packets until 70 d post explosion with logarithmically spaced time steps (). The initial time steps of the nebular phase simulation were also treated in LTE and the NLTE solver was switched on at time step 12 (28 d).
@@ 2.2 CO He WD merger ejecta model
| Element | Mass [] |
| He | |
| C | |
| O | |
| Ne | |
| Na | |
| Mg | |
| Al | |
| Si | |
| S | |
| Ar | |
| Ca | |
| Ti | |
| Cr | |
| Fe | |
| Co | |
| Ni | |
| 48Cr | |
| 52Fe | |
| 56Ni |
While artis is capable of multi-dimensional simulations, owing to the increased computational cost of including our full NLTE treatment, we work in 1D for this first exploratory study and employ a spherically averaged model. The full 3D model has clear asymmetries (see Morán-Fraile et al. 2024 or Kozyreva et al. 2024) which likely impact the strengths and shapes of spectral features. However, the 1D average preserves the main characteristics of the model which allows us to address the key questions of whether He i and [Ca ii] line formation are plausible. We construct our 1D model by spherically averaging the 3D hydrodynamic simulation into 87 radial bins. The composition of this 1D ejecta model is presented in Table LABEL:tab:ejecta_composition and shown in Figure 1 along with the energy deposition profiles at the epochs we present spectra for. From Figure 1 we can see that because the He distribution extends to low velocities there is significant overlap between where the He is present in the ejecta and where the deposition is taking place.
@@ 3 Results
Figures 2 and 3 show the optical and NIR spectra predicted by our simulations. We display three epochs with the r-max and +20 d r-max spectra taken from our photospheric phase simulation and the +45 d r-max spectrum taken from our nebular phase simulation444The spectra presented here will be made available on the Heidelberg supernova model archive HESMA (Kromer et al. 2017, https://hesma.h-its.org)..
@@ 3.1 Optical spectroscopic evolution
Ti ii remains a key species in the spectrum, contributing significantly across a range of wavelengths spanning the entire optical spectrum. There is a significant increase in the contributions of neutral species such as Fe i, Ca i, Cr i, Si i and Co i. This is a result of the reduced ionization state of the simulation at this epoch. In particular, Fe i has a substantial contribution across a wide range of optical wavelengths and the Fe i feature predicted at Å, which also has contributions from Co i and Ni i, is the most prominent feature in the spectrum other than the Ca ii/[Ca ii] features discussed above. The complex of emission features predicted by the simulation between - Å is also primarily attributed to the blended contributions of neutral species (Fe i, Ca i, Cr i, Co i).
@@ 3.2 NIR spectroscopic evolution
The most prominent feature in the peak NIR spectrum predicted by our simulation is the strong He i 10830 Å absorption feature (see Figure 3), while there is also a clear He i 20587 Å absorption feature. We note that substantial ejected He masses, such as the ejected for our model, are likely required for a He i 20587 Å feature to appear: previous artis simulations of CO WD explosions, ejecting of He, predicted clear He i 10830 Å features, but no obvious He i 20587 Å feature formed (Collins et al. 2023; Callan et al. 2025). A prominent Ca ii 11839, 11950 Å absorption feature is also present at Å. The NIR spectrum is dominated by Ti ii emission for wavelengths shorter than 16000 Å while redward of this He i provides the greatest contribution to the spectrum.
@@ 4.2 Model comparisons with observations in nebular phase
As can be seen from Figure 5 the nebular phase spectrum predicted by our simulation is dominated by the strong [Ca ii] and Ca ii emission features at around 7300 and 8700Å respectively. The simulated features have comparable strengths and velocities to the features observed in the nebular phase spectra of Ca-rich transients. However, our simulation predicts the Ca ii NIR triplet emission to be stronger than the [Ca ii] 7291, 7324 Å emission while these features have roughly equivalent strengths in the observed spectra555Radiative excitation still plays a role in the simulation at this time which keeps the NIR triplet relatively strong.. Additionally, while we note the observed Ca ii emission features show clear variations in their widths and velocities between different Ca-rich transients (see Figure 5), our simulation predicts Ca ii features that are too broad relative to the observations.
@@
Blueward of 7000 Å the observed Ca-rich transients show very substantial variation in their nebular phase spectra and as such the model has varying success in matching the observed spectra over this wavelength range. One feature that appears to be common to all the observations that we compare with here is the emission feature around 5500 Å, which the model attributes primarily to Fe i emission. Other than SN 2022oqm, all the Ca-rich transients which we compare to predict an emission feature around 6300 Å attributed to [O i] 6300, 6364 Å. However, while the simulation predicts such an emission feature the contribution of [O i] 6300, 6364 Å is quite minimal with blended emission from a variety of species including Fe i, Cr i, Ca i, Co i driving the formation of the feature (see Figure 2). While this suggests the [O i] 6300, 6364 Å feature is not a clean diagnostic at this epoch it will become more reliable at later times as forbidden emission increasingly dominates over the contribution of allowed transitions to the spectrum.
@@ 5 Conclusions
Therefore we conclude that low-mass CO WDHe WD mergers are promising candidates for Ca-rich transients. However, it is not yet clear whether such models can account for the full range of observed objects in this class of transients. Our model shows a strong spectral contribution from Ti ii at all epochs resulting in an SED at peak that is substantially redder than the majority of Ca-rich transients. The strong [Ca ii]/Ca ii emission features that dominate our simulated nebular spectra are too broad relative to observations and the He i 6678, 7065 Å absorption features predicted at peak appear to be more prominent than in the majority of Ca-rich transients. Interestingly the simulations presented by Zenati et al. (2023) for a CO WD disrupted by a hybrid HeCO WD also produce nebular spectra with strong [Ca ii] but in contrast do not display any strong contributions from He. Therefore, taken together these different WD mergers can produce a family of transients that are unified by strong [Ca ii] in their nebular spectra but show substantial variation in the strength of He spectral features predicted.
Future work should investigate whether the discrepancies between our model and observations can be resolved. In particular, the He and Ti distributions vary significantly with line-of-sight in the 3D hydrodynamic merger simulation, suggesting that observer orientation may be important. The strength of He and Ti features can also be impacted by factors relevant to their ionization state such as the distribution of in the ejecta which also shows significant variation with viewing angle in the 3D ejecta model. Additionally, the [Ca ii] feature can show asymmetric profiles and in some cases blue-shifts for observed Ca-rich transients, which may indicate that a multi-dimensional structure is present. Future multi-dimensional NLTE radiative transfer simulations should therefore be carried out to investigate if variation between lines-of-sight in the 3D ejecta model can better capture the variety observed for the Ca-rich class.
Ca-rich transients show a variation of approximately three magnitudes in peak r-band brightness. Therefore, further hydrodynamic merger simulations should be carried out to explore if systems with different combinations of CO and He WD masses can robustly explode resulting in a diversity in peak brightness which can cover the variation observed for the Ca-rich class. Additionally, if systems with a lower mass He WD can explode, this should result in a reduction in the He and Ti ejected in the explosion. Future simultaneous optical and NIR observations (e.g. from SOXS Schipani et al. 2018) will be ideal for the testing of such models by capturing both the overall SED and the key He i NIR features.
@@ Acknowledgements.
FPC and SAS would like to acknowledge support from the UK Science and Technology Facilities Council (STFC, grant number ST/X00094X/1). This work used the DiRAC Memory Intensive service (Cosma8) at Durham University, managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Durham was funded by BEIS, UKRI and STFC capital funding, Durham University and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). AH is a fellow of the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD) and acknowledges financial support from IMPRS-HD. This work has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101152610. LJS acknowledges support by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (ERC Advanced Grant KILONOVA No. 885281). JMP acknowledges the support of the Department for Economy (DfE). This work received support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 759253 and 945806, the Klaus Tschira Foundation, and the High Performance and Cloud Computing Group at the Zentrum für Datenverarbeitung of the University of Tübingen, the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant no INST 37/935-1 FUGG. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – RO 3676/7-1, project number 537700965, and by the European Union (ERC, ExCEED, project number 101096243). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. SAS thanks the Cosmic Dawn Center / Niels Bohr Institute (University of Copenhagen) who hosted his sabbatical visit during which part of this work was carried out. NumPy and SciPy (Oliphant 2007), Matplotlib (Hunter 2007) and artistools666https://github.com/artis-mcrt/artistools/ (Shingles et al. 2025) were used for data processing and plotting.References
@@
- Callan et al. (2025) Callan, F. P., Collins, C. E., Sim, S. A., et al. 2025, MNRAS, 539, 1404
@@
- Rosswog et al. (2009) Rosswog, S., Ramirez-Ruiz, E., & Hix, W. R. 2009, ApJ, 695, 404
- Schipani et al. (2018) Schipani, P., Campana, S., Claudi, R., et al. 2018, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 10702, Ground-based and Airborne Instrumentation for Astronomy VII, ed. C. J. Evans, L. Simard, & H. Takami, 107020F
@@
- Sell et al. (2015) Sell, P. H., Maccarone, T. J., Kotak, R., Knigge, C., & Sand, D. J. 2015, MNRAS, 450, 4198
- Shen et al. (2010) Shen, K. J., Kasen, D., Weinberg, N. N., Bildsten, L., & Scannapieco, E. 2010, ApJ, 715, 767